Recent Log Entries

Results of Hall B BLA IBC2C24 signal injection

Log Number: 3829920Logbook(s): Body: 

Fig 1: The final hour of data in IBC2C24, prior to the shutdown for the SAD. Noise is evident.

Fig 2: Signal injection on the test port, representing various beam currents. No noise evident, especially not in comparison to Fig 1.

Fig 3: Confidence test results: 1195 nA / 0.24 (cal factor) = 4.98 uA.

---------

This test rules out the cavity cable/connectors, and electronics, as a source of the noise in question.

Vertical Separator Pass 5 SSA 3 Tripped

Log Number: 3829916Logbook(s): Entry Makers: wissmannBody: 

While working in building 68 I noticed the back room that houses the Vertical ssAs was very warm - the AC wasn't cooling. Also noticed Cavity 3's SSA had tripped.
Archived forward power showed the trip happened around 7:05. The SSA tripped and showed a mismatch and DC power fault - see the attached figure. Notified the
Crew Chief about the AC and SSA trip and started an SSA restart.

heat vs trips NL

Log Number: 3829909Logbook(s): Entry Makers: beneschBody: 

1L05-8 produces too much heat for usual circumstances. Others in the zone aren't much better, which is why it is getting pulled in October and C75-1 installed. Figure 2 shows that we have a fair amount of heater margin in the NL. lem predicts that NL fault rate will drop 3/shift is 1L05-8 is brought back up at 3 MV/m. If CAPHTRMGN doesn't drop by more than 3 W (*18) at this setting, go ahead.

Autolog save

Log Number: 3829905Logbook(s): Tags: Body: saver ran from opsl632

Saved: ywang 08-18-20 11:02:17 {A:1 B:5 C:5 E:1050 A=Tune B=BS C=BS D=BS 0R08=0.00 0R08 dump in. A B C dump(s) in.

Inj work summary

Log Number: 3829904Logbook(s): Entry Makers: ywangBody: 

Optimization of injector:

1. Adjusted orbit in 0I and 01 regions
2. Restored HWien setting
3. Increased prebuncher amplitude from 8 to 9
4. Backed out RF 0L04 gang phase offset by 1 deg
5. Verified no leak from A or C to B at chopper
6. Hall A verified PQB

FaultViewer

Log Number: 3829903Logbook(s): Entry Makers: beneschBody: 

I updated the ODVH on the top two offenders. It is hoped that the SL repairs reduce the trips in the rest of the energy lock zone. I see on figure 3 that 2L26-2,4 are back up.

IHAGT06A - grounding cable gauge changed

Log Number: 3829899Logbook(s): Entry Makers: delacruzBody: 

Verified the harp assembly was grounded to the building ground. Used a heavier gauge grounding cable and connected it from the preamp to the building ground. The ADC offset is still around 92mV.

Re-check Hall B noise after Injector steering

Log Number: 3829894Logbook(s): Entry Makers: suleimanBody: 

Beam motion noise is still there (as expected). The noise on Hall B beam current is a result of this motion on tight appretures especially when we have bad transmission.

No leak from A or C to B

Log Number: 3829882Logbook(s): Entry Makers: ywangBody: 

A and C slits closed, B slit open to 17.65:

Fig. 1: All beam off, Fcup# = 0.0106
Fig. 2: A at high current setting (135uA), Fcup# = 0.0106
Fig. 3: C at high current setting (40uA), Fcup# = 0.0106

It does not appear to be any leakage from A or C to B.

Pages